Suppose a researcher goes to a small college of 200 faculty, 12 of which have blood type O-negative. He obtains a simple random sample of n = 20 of the faculty. Let the random variable X represent the number of faculty in the sample of size that have blood type O-negative.

(a) What is the probability that 3 of the faculty have blood type O-negative? (Round your answer to the nearest thousandth).


1
Expert's answer
2021-06-07T11:13:59-0400

"n=20 \\

p=\frac{12}{200}=0.06 \\

q=1-0.06=0.94 \\

P(X=3) = \frac{20!}{3!(20-3)!}(0.06)^3(0.94)^{20-3} \\

= \frac{6840}{6} \times 0.000216 \times 0.349279 \\

=0.086"